OpenStudy (anonymous):

solve the silmutaneous question 3x+y=2 and 3x^2 +y^2+xy=6

OpenStudy (anonymous):

Using substitution method, 3x+y = 2 y=2-3x \[3x^2+(2-3x)^2+x(2-3x)=6\] \[3x^2+4-12x+9x^2+2x-3x^2-6=0\] \[9x^2-10x-2=0\] x=1.284 or -0.173